Log record of  recent activity and
status update of cameras
17. October 2014
Installation of SOAO camera system.
Camera operation is set to: Start: 6pm. Stop: 6am.
17. November 2014
Installation of BOAO camera system.
Camera operation is set to: Start: 6pm. Stop: 6am.
25. Decmber 2014
Failure of SOAO Cam1
(test were carried out with the help of SOAO staff members. The problem must be with the camera itself). Repair is scheduled in Q1 of 2015
9. January 2015
Semi-failure of SOAO Cam3
(the camera is not automatically recognised by UFOCaptureV2. It has to be manually enabled within the software. This is a minor problem as it can be initiated manually via VNC login).
28. January 2015
Received picture from Mr. Cho telling and showing that BOAO camera system looks good. Outside temperatures are below freezing point.
07. March 2015
BOAO Meteor PC down after Win7 update. There seem to be a problem with rebooting properly.
08. March 2015
BOAO Meteor PC down after Win7 update. There seem to be a problem with rebooting properly.
09. March 2015
BOAO Meteor PC is up and running again.
16. March 2015
Interesting sky activity on SOAO Cam1 & Cam3 - check out videos!
17. March - 02. April 2015
BOAO Cam3 was down, now up and running again from 02. April 2015
26. March 2015
Night with many triggers due to interference noise. Not sure why. But the resulting JPG images can be used to create a time-laps GIF movie.
03. April 2015
Inspection of SOAO meteor cameras. All cameras were checked. All cams are functioning. That means they survived their first 2014/2015 winter season.  The CCTV camera housing heater must have worked. The reason for failure of SOAO Cam1 was probed and identified. The fault is a defect underground coaxial cable ofCam1. Most likely due to mouse bite attack. Also I noticed that power connections in the power/fuse/timer plastic box is not optimal. Also I identified the semi failure (see 09. January 2015). The reason for it was a bad connection of the A/D converter in the wrong USB port. Now it seems to work again and the camera software (UFOCaptureV2) start automatically.
03. April 2015
Also: SOAO Cam1 and SOAO Cam3 were physically removed from their respective CCTV housing (as part of testing for their functionality). This will have a change in the astrometric solution for detections obtained after 03. April 2015 as the orientation (mostly camera rotation) might have changed a little.
03. April 2015
Maintenance work: Cleaned all frontglass windows of CCTV camera houses. Also: oiled (with WD40) all screws and locks.
06. April 2015
Power ON/OFF timer at SOAO was changed. Start: 8pm. Stop: 5am.
09. April 2015
Power ON/OFF timer at BOAO was changed. Start: 8pm. Stop: 5am.
Month of May 2015
SOAO received an upgraded powergenerator which resulted in several power- and internet-outages. This affected the normal operation of the meteor cameras and the PCs were shut-down several times. Also I noticed that the dimension4 time synchronisation software did not synchronise as usually in previous months. This is most likely because of failed internet connection.
Also during May the three SOAO front glass windows were cleaned from asian yellow dust.
Around 15. June 2015
SOAO Observatory was struck by a lightening. Possibly the SOAO Meteor PC is damaged.
08. July 2015
Visited SOAO and briefly inspected the SOAO camera system after lightening strike (causing an EMP - electromagnetic pulse) which can be confirmed by inspecting the Hauppage Impact VCB PCIe video cards. One of them showed clear signs of burning at the composite connectors. Both VCB cards are recognised in the device manager and flagged as "working". But no video signal is received when testing with UFOCapture and Win Movie Maker. The EzCap A/D converter does not work either. Also the SOAO Meteor PC is gone (does not boot upon pressing power button). But both 2GB RAM modules work and the HDD works so I could copy files to the new computer (generously sponsored by KASI). However, since there was an SPD (Surge Protection Device) infront of each camera, they survived!! Lucky me!! I tested all cameras by physically removing them from the CCTV housing and take them inside for inspection. They all work fine. Also the over- and underground coax cables were tested. They all work fine. Especially Camera-1 cable is working fine. There was no damage due to mouse bite. What happened here is that I might by mistake have turned the camera IRIS knob into the wrong direction so that it closes the iris or wrong software setting in UFOCapture (set to PAL instead of NTSC in the video settings). That of course gives you a black screen and gives the impression that it does not work. For the future I think the following is necessary: Either install an SPD also infront of each video A/D card so to protect the computer or pull an optical cable and use multiplexers for data transmission. Optical cables are immune against EMPs. If I chose the first solution, I think the system should be shut-down (all cables disconnected) during the thunderstorm/rainy season here in Korea which is between May and September. Lets see.
01. - 04. October 2015
Visiting SOAO: Restarting SOAO meteor cams on 01. October 2015. New A/D video converters (Pinnacle Dazzle DVC100) were installed.
Timer/Capture start: 6pm
Timer/Capture stop: 6am
All camera housing front glass cleaned.
BIAS frames (useful to identify dead pixels) for each camera (short video sequence) were taken on a cloudy night on 01. October 2015.
HOWEVER: NO SPDs ARE CURRENTLY INSTALLED.
25. April 2016
Visiting SOAO: After months of trial and various tests at SOAO I finally figured out that connecting two cameras on the same USB hub (two USB connectors being connected on the same place on the main board) is not a good idea. It causes the Windows machine to crash with a Blue Screen of Death (BSoD) reboot. Learning lesson: One camera, one USB connector.
29. May 2016
Visiting SOAO: All cameras switched off (main power supply: OFF) and coaxial cables (both ends) disconnected due to danger of lightening strike. SOAO Meteor PC was also switched off. I verified the following:
Camera 1 --> Video: 3 Dazzle
Camera 2 --> Video: 1 Dazzle
Camera 3 --> Video: 2 Dazzle
The cameras will be back online around August/September 2016.
14. June 2016
Visiting BOAO: Detected a faulty 12V camera powersupply. New one will be purchased soon for installation/repairment in Aug/Sep 2016. To protect against lightening during the Korean raining season (June/July/Aug) strikes the power and coaxial cables (both at the camera and PC end) has been switched off and disconnected. Before that all cameras and connections were tested and are functioning in good order. I verified the following:
Camera 1 --> Video: 1
Camera 2 --> Video: 2
Camera 3 --> Video: 3
The BOAO Meteor PC has been switched off. No BOAO observations during summer 2016.
26. September 2016
Visiting BOAO: Replaced faulty 12V powersupply. Camera power supply was switched on. Video cables were reconnected without surgeprotection. Cameras were tested and all worked well. Front glasses for each camera housing were cleaned.
17. - 20. October 2016
Visiting SOAO: Replaced old HDD with a new 1GB HDD. Tested the new drive for errors. All OK. Installed Win 7 from scratch. Installed all newest updates and various software and drivers. I verified the following:
Camera 1 --> Video: 1 (UFOCapV2)
Camera 2 --> Video: 2 (UFOCapV2)
Camera 3 --> Video: 3 (UFOCapV2)
Also I filled silicone in cables holes in the gray (video & electric) boxes to prevent bugs and insects to enter the boxes via the cable protection pipes. Cameras were tested and work well. Front glass were cleaned. SOAO Meteor Detection PC seems to work stable.
14. May 2017
BOAO: BOAO is off-line. Power is OFF. Disconnected all three coaxial cables on both ends (camera and pc)
to prevent damage due to lightening strike.
18. May 2017
SOAO: SOAO is off-line. Power is OFF. Disconnected all three coaxial cables on both ends (camera and pc)
to prevent damage due to lightening strike. Also USB cables were unplugged on PC end by Taek-su.